# Funding AI for Good: A Call for Meaningful Engagement

### v2

@@ 2.2.1. Techno-Centric Approaches
A techno-centric approach, as defined by  Dodson et al. (2012), “place[s] the technology at the center of the intervention by giving prominence to technical features and capabilities.” It is rooted in “techno-solutionism,” where complex, structural challenges are reduced to problems amenable to technical interventions (Morozov, 2013), often under the false assumption that these solutions are neutral or apolitical (Feenberg, 2002; Winner, 2010). While leveraging emerging technology as a silver bullet to social problems appears optimistic and well-intentioned, Berlant’s concept of “cruel optimism” points to the cruel side of that optimism: promising social transformation while deepening surveillance, reproducing bias, or ignoring community needs (Berlant, 2011). These initiatives often fail due to “design-reality gaps” (Heeks, 2002): a mismatch between system designers’ perceptions of local contexts and local actuality.
In “tech for good”, techno-centric projects are driven by an “innovate at all costs” mentality, leading to the creation of technologies with little maintenance support or local capacity building afterward (Saha et al., 2022b). It incentivizes a top-down design approach, where the solutions are typically designed by external technologists without adequate engagement with the local communities (Walsham, 2017; Dada, 2006; Harris, 2016; Dodson et al., 2012), especially during the decision-making stages of projects (Saha et al., 2022a). It could also lead to the phenomenon of “bungee jumping research” (Dearden and Tucker, 2016), characterized by brief, superficial engagement (Saha et al., 2022b; Sultana et al., 2019; Dada, 2006). Development of the solutions tends to follow a linear and short-term process: building the technology, deploying it, and moving on to the next project, where researchers often leave after their work is published, making it difficult to sustain the systems they have introduced (Taylor et al., 2013; Dada, 2006; Walsham, 2017). Last but not least, evaluations of techno-centric projects constrain assessments of success to technical dimensions, narrowly focusing on metrics such as model accuracy or the production of tangible artifacts like software (Dodson et al., 2012; Lin et al., 2024).
@@ 2.2.2. Balanced Approaches
A balanced approach is grounded in the understanding of problems faced by specific communities, in addition to leveraging emerging technology’s capacities. Community-collaborative approaches (CCA) are central to a balanced approach. A systematic review by Cooper et al. defines CCA as projects that “involve collaboration with community stakeholders as co-researchers throughout the research process while investigating complex problems of concern to the community (e.g., social and health-related), and developing novel technologies or sociotechnical solutions” (Cooper et al., 2022). CCA includes work in participatory design (e.g., (Delgado et al., 2023; Bondi et al., 2021)), community-based participatory research (e.g. (Wan et al., 2023; Liang et al., 2023; Strohmayer et al., 2017)), and action research (e.g., (Hayes, 2011; Le Dantec, 2016; Lemaire and Muñiz, 2011; Balestrini et al., 2014)). Projects that follow CCA are more likely to create lasting social impact because they have community support and input from the start (Gandhi et al., 2007) and amplify positive values already present in the community (Toyama, 2015)Brown and Mickelson (2019) emphasize the necessity of “the full inclusion of all relevant parties in every aspect of the project” to ensure project success.
A balanced approach emphasizes that community members’ lived experiences are a legitimate form of knowledge within their specific contexts (Erete et al., 2023; Cooper et al., 2022), on par with technical expertise. This emphasis is grounded in the concept of epistemic justice (Fricker, 2007), which calls for fairness in recognizing who is seen as a knower and what forms of knowledge are valued. In “tech for good” projects, this means embracing modalities of knowledge sharing that go beyond traditional academic outputs, such as storytelling and community observations (Erete et al., 2023; Harrington et al., 2019a; Zegura et al., 2018)Zegura et al. (2018) demonstrate that valuing the community’s local knowledge is not only crucial for the quality of data and the subsequent data analysis, but also generates “care” among researchers and community collaborators (Zegura et al., 2018). Relationships among collaborative teams are critical for the feasibility and sustainability of community-based projects (Dell and Kumar, 2016; Kumar and Dell, 2018). Sustainable relationships with communities can also be cultivated through regular presence, trust-building, and iterative design processes, and they enable project agendas to emerge organically from (rather than being imposed on) communities (Dearden and Tucker, 2016; Carroll and Rosson, 2013). Lastly, a balanced approach extends evaluation beyond technical outputs to include community outcomes such as increased technical capacity (Zegura et al., 2018), impact on everyday practice (Taylor et al., 2016), and “empowerment of the participants and the knowledge gained by the technical designers” (Drain et al., 2021).
@@
Prior work has also highlighted some limitations of CCA (Pine et al., 2020; Erete et al., 2023; Harrington et al., 2019b). For example, nonconsensual, short-term, and non-contextual-specific participation would lead to the exploitation of impacted communities through “participation washing” (Sloane et al., 2020). Sloane et al. argue for recognition of participation as work that is valuable for machine learning development, including data work and context-specific consultation (Sloane et al., 2020). In AI4SG specifically, Lin et al. document the contributions that community organization staff make to AI4SG projects, including data work and participating in regular feedback meetings (Lin et al., 2024). However, this labor is often invisible or overlooked by downstream AI developers and the public (Crain et al., 2016; Gray and Suri, 2019; Crawford, 2021; Sambasivan et al., 2021). A recent body of work has started to quantify invisible labor in crowd work (e.g., managing payments and hyper-vigilance) (Toxtli et al., 2021) and online volunteer moderators (e.g., approving content and managing users) (Li et al., 2022), revealing that such unseen tasks make up a substantial portion of the overall work.
@@ 2.3. Emerging Understanding of “Tech for Good” Decision-Makers and AI4SG Investments
Focusing on private tech companies specifically, some recent works have started to critically examine investment in AI4SG through case study analysis (Nost and Colven, 2022; Lukacz, 2024; Henriksen and Richey, 2022). For example, Nost and Colven (2022) view Microsoft’s AI for Earth program as a form of greenwashing and “philanthrocapitalism”, where the company uses the initiative to enhance its reputation as a socially responsible corporation, while simultaneously contributing to environmental harm through its broader business operations, including its reliance on fossil fuels for data centers. Henriksen and Richey (2022) examine Google’s philanthropic efforts, similarly arguing that they serve dual purposes: advancing social causes while simultaneously expanding Google’s market influence and integrating its technologies into more areas of life. Comparing “Big Tech” companies with the “Big Tobacco” industry, Abdalla and Abdalla (2021) examined “Big Tech” funding’s influence on the broader academic landscape (though not specific to AI4SG), shaping events and research questions that “put forward a socially responsible public image” of tech companies.
@@ 2.4. Document Analysis of Pre-Existing Texts about AI and Funding (beyond AI)
HCI scholars have used document analysis to examine pre-existing texts about AI, encompassing AI ethics toolkits (Wong et al., 2023b), AI policy documents (Wong et al., 2023a; Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), AI risk manifesto-style statements (Oldenburg and Papyshev, 2025), patents (Cheon, 2023), and military funding (Widder et al., 2024). Pre-existing documents about AI do not simply describe AI as facts; they construct AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025), rooted in Jasanoff and Kim’s “sociotechnical imaginaries” (collective visions of technology’s role in society) (Jasanoff and Kim, 2009). These documents not only define what AI is, what problems it should solve, who benefits from it, and whose perspectives remain invisible, they also “actively shape the trajectory of AI development, integration, and governance” (Zhong et al., 2025). For example, Wong et al. (2023b) analyzed a corpus of 27 AI ethics toolkits to “identify the discourses about ethics, the imagined users of the toolkits, and the work practices the toolkits envision and support.” They found that AI is imagined as a technology that can be made “ethical” through technical adjustments—better models, checklists, or audits (Wong et al., 2023b)Sophie Bennani-Taylor (2024) argue that policy documents perform “discursive infrastructuring”: they stabilize certain meanings of AI and legitimize particular sociotechnical arrangements. Through an analysis of UK’s National AI Strategy document, the author found that AI is constructed as 1) an autonomous, inevitable force, 2) a driver of national progress and economic modernization, and 3) a domain dominated by technical expertise over social knowledge (Sophie Bennani-Taylor, 2024). Closest to the corpus analyzed in our work, Widder et al. (2024) analyzed U.S. Department of Defense’s grant solicitations for the use of AI in military applications and revealed a deep entanglement between academic AI research and militarized imaginaries, where recurring rhetorical devices like “one small problem” justify continuous funding increases for military AI.
@@
Similarly, prior research examining funding documents (not limited to AI) shows that funding languages are not neutral; they encode priorities, shape applicant behaviors, and structure downstream research collaborations. For example, Palmero et al. (2025) conducted a textual network analysis of Horizon Europe funding calls, treating these texts as “cognitive artefacts of policy design” that reveal how high-level policies are translated into operational agendas. The authors point out that funding document “predefines applicants’ behaviour, shapes their own perceptions of eligibility, and influences evaluative processes” (Palmero et al., 2025). Existing work in innovation studies likewise demonstrates the downdream influence of funding documents: Peng et al. (2024) find that the proportion of ”promotional language” (such as “innovative” and “unique”) in grant proposals is statistically associated with a proposal’s likelihood of being funded, while Smith et al. (2023) show “how thematic directives rooted in funding calls influenced collaboration across research.”
Building on this body of work, we contribute a first document analysis of AI4SG funding documents to understand how funding language discursively constructs AI imaginaries (especially AI’s role in social good domains) and characterizes the conditions under which its positive impact might be achieved.
@@ 3.1. Qualitative Dataset Development
In this paper, we define “funders” as organizations that provide monetary support to AI4SG, such as philanthropic foundations, nonprofit organizations, government agencies, and private technology companies. While individuals may also fund AI4SG projects, we focused on organizations because they are more likely to provide public articulations of their funding agendas. We define “funding programs” as specific AI4SG initiatives, such as a research grant on the intersection of climate change and AI. We define “funding documents” as funding calls for proposals and grant announcements through which funders communicate their intentions and envisioned approaches in AI4SG. While funding calls typically outline more detailed guidelines, grant announcements present a more concise overview of an initiative’s aims, scope, and thematic priorities. We consider both as funding documents because they offer complementary insights into funding agendas.
@@
We conducted the search with keywords indicating 1) the emphasis on social good (“social good,” “social impact,” “climate change,” “public health,” “justice,” “equity,” and “community”), 2) the methodological orientation (“Artificial Intelligence,” “AI,” “Machine Learning,” “ML,” or “Data Science”), and 3) the focus on funding (“funding,” “fund,” “award,” or “grant”). The query returned 296 links to web pages. The links led to calls for proposals, announcements about the grants, blog posts about AI4SG projects, academic publications in AI4SG, academic departments and labs working on AI4SG, and so on. Using snowball sampling (Naderifar et al., 2017), we followed references within the initial links to identify additional funding programs. For example, a university funding opportunities page [^1] https://e-hail.umich.edu/funding-opportunities/ accessed Aug 2, 2025 listed multiple AI4SG-related programs not captured in the original search. We additionally examined these funding programs for inclusion.
@@
Using these criteria, the first author then examined the remaining links, and the second author reviewed the selections. The first two authors then resolved any disagreements. We narrowed down to 35 funding documents (19 funding calls and 16 grant announcements) for in-depth qualitative analysis. The funding programs represent 23 funders and around \$410 million in investments in aggregate. We also included supplementary materials referenced in funding documents in our qualitative dataset, as they offer important details about the funding programs, such as Frequently Asked Questions and application forms. For example, we supplemented the Google AI Impact Challenge funding call (Google, 2018) with its application guide (Google, 2019), which laid out selection criteria and application questions.
@@
Table LABEL:tab:funding_doc_mr in Appendix A provides the list of funding documents and information about the funding programs (e.g., funding amount, duration, supporting funders, target regions). Our dataset covers a wide range of programs from 2017 to 2025, aiming at both real-world implementation and use-inspired research. These funding documents were provided by government agencies (11 funding documents), philanthropic foundations (10), nonprofit organizations (7), and private technology companies (8) based in English-speaking countries. These funders were headquartered in the United States (26), the United Kingdom (5), Canada (4), Sweden (3), Germany (1), Ireland (1), and Australia (1). These funding programs targeted applications from countries including the United States (5), the United Kingdom (1), Ireland (1), and South Africa (1), as well as broader regions including globally (14), Africa (9), Latin America (2), and Low- and Medium-Income Countries (2).
@@ 3.2. Thematic Analysis on Funding Documents
In the third coding stage, the first author coded all the funding documents, with our research questions in mind. Each funding document’s codes were reviewed by another author in our team. The first author then resolved any (though few) disagreements (denoted as comments on Google documents) with each code reviewer through online discussions. These discussions involved clarifying the meaning of the first author’s original codes and discussing whether they aligned with the original texts in the funding documents. We did not calculate intercoder reliability, since reflexive thematic analysis treats coding as an interpretive process rather than one that seeks positivist (quantitative) reliability measures (Braun and Clarke, 2019). All authors then conducted iterative affinity diagramming on an online whiteboard and discussed emerging high-level themes from the codes. The purpose of affinity diagramming was to categorize low-level codes (in the sentence level) and surface higher-level themes across the funding documents (in the section level and beyond), as well as finding connections between these themes. We collaboratively characterized these high-level themes into a spectrum between a more techno-centric approach or a more balanced approach using the lens of prior literature synthesized in Section 2.2. Appendix B provides screenshots of our coding process and affinity diagrams.

### v3

@@ 2.2.1. Techno-Centric Approaches
A techno-centric approach, as defined by Dodson et al. (2012), “place[s] the technology at the center of the intervention by giving prominence to technical features and capabilities.” It is rooted in “techno-solutionism,” where complex, structural challenges are reduced to problems amenable to technical interventions (Morozov, 2013), often under the false assumption that these solutions are neutral or apolitical (Feenberg, 2002; Winner, 1980). While leveraging emerging technology as a silver bullet to social problems appears optimistic and well-intentioned, Berlant’s concept of “cruel optimism” points to the dark side of that optimism: promising social transformation while deepening surveillance, reproducing bias, or ignoring community needs (Berlant, 2011). These initiatives often fail due to “design-reality gaps” (Heeks, 2002): a mismatch between system designers’ perceptions of local contexts and local actuality.
In “tech for good”, techno-centric projects are driven by an “innovate at all costs” mentality, leading to the creation of technologies with little maintenance support or local capacity building afterward (Saha et al., 2022b). This mentality incentivizes a top-down design approach, where the solutions are typically designed by external technologists without adequate engagement with the local communities (Walsham, 2017; Dada, 2006; Harris, 2016; Dodson et al., 2012), especially during the decision-making stages of projects (Saha et al., 2022a). It could also lead to the phenomenon of “bungee jumping research” (Dearden and Tucker, 2016), characterized by brief, superficial engagement (Saha et al., 2022b; Sultana et al., 2019; Dada, 2006). Development of the solutions tends to follow a linear and short-term process: building the technology, deploying it, and moving on to the next project. Researchers often leave after their work is published, making it difficult to sustain the systems they have introduced (Taylor et al., 2013; Dada, 2006; Walsham, 2017). Last but not least, evaluations of techno-centric projects constrain assessments of success to technical dimensions, narrowly focusing on metrics such as model accuracy or the production of specific artifacts like software (Dodson et al., 2012; Lin et al., 2024).
@@ 2.2.2. Balanced Approaches
A balanced approach is grounded in the understanding of problems faced by specific communities, in addition to leveraging emerging technology’s capacities. Community collaborative approaches (CCA) are central to a balanced approach. A systematic review by Cooper et al. (2022) defines CCA as projects that “involve collaboration with community stakeholders as co-researchers throughout the research process while investigating complex problems of concern to the community (e.g., social and health-related), and developing novel technologies or sociotechnical solutions”. CCA includes work in participatory design (e.g., (Delgado et al., 2023; Bondi et al., 2021)), community-based participatory research (e.g. (Wan et al., 2023; Liang et al., 2023; Strohmayer et al., 2017)), and action research (e.g., (Hayes, 2011; Le Dantec, 2016; Lemaire and Muñiz, 2011; Balestrini et al., 2014)). Projects that follow CCA are more likely to create lasting social impact because they have community support and input from the start (Gandhi et al., 2007) and because they amplify positive processes already present in the community (Toyama, 2015). Brown and Mickelson (2019) emphasize the necessity of “the full inclusion of all relevant parties in every aspect of the project” to ensure project success.
A balanced approach emphasizes that community members’ lived experiences are a legitimate form of knowledge within their specific contexts (Erete et al., 2023; Cooper et al., 2022), on par with technical expertise. This emphasis is grounded in the concept of epistemic justice (Fricker, 2007), which calls for fairness in recognizing who is seen as a knower and what forms of knowledge are valued. In “tech for good” projects, this means embracing modalities of knowledge sharing that go beyond traditional academic outputs, such as storytelling and community observations (Erete et al., 2023; Harrington et al., 2019a; Zegura et al., 2018). Zegura et al. (2018) demonstrate that valuing the community’s local knowledge is not only crucial for the quality of data and the subsequent data analysis, but also generates care among researchers and community collaborators (Zegura et al., 2018). Relationships among collaborative teams are critical for the feasibility and sustainability of community-based projects (Dell and Kumar, 2016; Kumar and Dell, 2018). Sustainable relationships with communities can be cultivated through regular presence, trust-building, and iterative design processes, and they enable project agendas to emerge organically from (rather than being imposed on) communities (Dearden and Tucker, 2016; Carroll and Rosson, 2013). Lastly, a balanced approach extends evaluation beyond technical outputs to include community outcomes such as increased technical capacity (Zegura et al., 2018), impact on everyday practice (Taylor et al., 2016), and “empowerment of the participants and the knowledge gained by the technical designers” (Drain et al., 2021).
@@
Prior work has also highlighted some limitations of CCA (Pine et al., 2020; Erete et al., 2023; Harrington et al., 2019b). For example, nonconsensual, short-term, and non-contextual-specific participation would lead to the exploitation of impacted communities through “participation washing” (Sloane et al., 2020). Sloane et al. (2020) argue for recognition of participation as work that is valuable for machine learning development, including data work and context-specific consultation. In AI4SG specifically, Lin et al. (2024) document the contributions that community organization staff make to AI4SG projects, including data work and participating in regular feedback meetings. However, this labor is often invisible or overlooked by downstream AI developers and the public (Crain et al., 2016; Gray and Suri, 2019; Crawford, 2021; Sambasivan et al., 2021).
@@ 2.3. Emerging Understanding of “Tech for Good” Decision-Makers and AI4SG Investments
Focusing specifically on private tech companies, recent work has begun to critically examine investment in AI4SG through case study analysis (Nost and Colven, 2022; Lukacz, 2024; Henriksen and Richey, 2022). For example, Nost and Colven (2022) view Microsoft’s AI for Earth program as a form of greenwashing and “philanthrocapitalism”, where the company uses the initiative to enhance its reputation as a socially responsible corporation, while simultaneously contributing to environmental harm through its broader business operations, including its reliance on fossil fuels for data centers. Henriksen and Richey (2022) examine Google’s philanthropic efforts, similarly arguing that they serve dual purposes: advancing social causes while simultaneously expanding Google’s market influence and integrating its technologies into more areas of life. Comparing “Big Tech” companies with the “Big Tobacco” industry, Abdalla and Abdalla (2021) examined the influence of “Big Tech” funding on the broader academic landscape (though not specifically on AI4SG), shaping events and research questions that “put forward a socially responsible public image” of tech companies.
@@ 2.4. Document Analysis of Pre-Existing Texts about AI and Funding (beyond AI)
HCI scholars have used document analysis to examine pre-existing texts about AI, encompassing AI ethics toolkits (Wong et al., 2023b), AI policy documents (Wong et al., 2023a; Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), AI risk manifesto-style statements (Oldenburg and Papyshev, 2025), patents (Cheon, 2023), and military funding (Widder et al., 2024). Pre-existing documents about AI do not simply describe AI as facts; they construct AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025), rooted in Jasanoff and Kim’s concept of sociotechnical imaginaries (collective visions of technology’s role in society) (Jasanoff and Kim, 2009). These documents not only define what AI is, what problems it should solve, and who benefits from it, they also “actively shape the trajectory of AI development, integration, and governance” (Zhong et al., 2025). For example, Wong et al. (2023b) analyzed a corpus of 27 AI ethics toolkits to “identify the discourses about ethics, the imagined users of the toolkits, and the work practices the toolkits envision and support.” They found that AI is imagined as a technology that can be made “ethical” through technical adjustments—better models, checklists, or audits (Wong et al., 2023b). Sophie Bennani-Taylor (2024) argue that policy documents perform “discursive infrastructuring”: they stabilize certain meanings of AI and legitimize particular sociotechnical arrangements. Through an analysis of UK’s National AI Strategy document, the author found that AI is constructed as 1) an autonomous, inevitable force, 2) a driver of national progress and economic modernization, and 3) a domain dominated by technical expertise over social knowledge (Sophie Bennani-Taylor, 2024). Closest to the corpus analyzed in our work, Widder et al. (2024) analyzed U.S. Department of Defense’s grant solicitations for the use of AI in military applications and revealed a deep entanglement between academic AI research and militarized imaginaries, where recurring rhetorical devices justify continuous funding increases for military AI.
@@
Similarly, prior research examining funding documents (not limited to AI) shows that funding languages are not neutral; they encode priorities, shape applicant behaviors, and structure downstream research collaborations. For example, Palmero et al. (2025) conducted a textual network analysis of Horizon Europe funding calls, treating these texts as “cognitive artifacts of policy design” that reveal how high-level policies are translated into operational agendas. The authors point out that funding document “predefines applicants’ behaviour, shapes their own perceptions of eligibility, and influences evaluative processes” (Palmero et al., 2025). Existing work in innovation studies likewise demonstrates the downstream influence of funding documents: Peng et al. (2024) find that the proportion of “promotional language” (such as “innovative” and “unique”) in grant proposals is statistically associated with a proposal’s likelihood of being funded, while Smith et al. (2023) show “how thematic directives rooted in funding calls influenced collaboration across research.”
@@ 3.1. Qualitative Dataset Development
In this paper, we define “funders” as organizations that provide financial support to AI4SG initiatives, including philanthropic foundations, nonprofit organizations, government agencies, and private technology companies. While individuals may also fund AI4SG projects, we focused on organizations because they are more likely to provide public texts about their funding agendas. We define “funding programs” as specific AI4SG initiatives, such as a research grant on the intersection of climate change and AI. We define “funding documents” as funding calls for proposals and grant announcements through which funders communicate their intentions and envisioned approaches in AI4SG. While funding calls typically outline more detailed guidelines, grant announcements present a more concise overview of an initiative’s aims, scope, and thematic priorities. We consider both as funding documents because they offer complementary insights into funding agendas.
@@
We conducted the search with keywords indicating 1) the emphasis on social good (“social good,” “social impact,” “climate change,” “public health,” “justice,” “equity,” and “community”), 2) the methodological orientation (“Artificial Intelligence,” “AI,” “Machine Learning,” “ML,” or “Data Science”), and 3) the focus on funding (“funding,” “fund,” “award,” or “grant”). The query returned 296 web page links. The links led to calls for proposals, announcements about the grants, blog posts about AI4SG projects, academic publications in AI4SG, academic departments and labs working on AI4SG, and so on. Using snowball sampling (Naderifar et al., 2017), we followed references within the initial links to identify additional funding programs. For example, a university funding opportunities page [^1] https://e-hail.umich.edu/funding-opportunities/ accessed Aug 2, 2025 listed multiple AI4SG-related programs not captured in the original search. We additionally examined these funding programs for inclusion.
@@
Using these criteria, the first author then examined the remaining links, and the second author reviewed the selections. The first two authors then resolved any disagreements. We narrowed down to 35 funding documents (19 funding calls and 16 grant announcements) for in-depth qualitative analysis. The funding programs represent 23 funders and aggregate investments of around \$410 million. We also included supplementary materials referenced in funding documents in our qualitative dataset, as they offer important details about the funding programs, such as Frequently Asked Questions and application forms. For example, we supplemented the Google AI Impact Challenge funding call (Google, 2018) with its application guide (Google, 2019), which laid out selection criteria and application questions.
@@
Table LABEL:tab:funding_doc_mr in Appendix A provides the list of funding documents and information about the funding programs (e.g., funding amount, duration, supporting funders, target regions). [^2] The full dataset can be accessed through this link:
https://github.com/nicole-hjlin/chi2026-ai-for-good-funding-dataset.
Our dataset covers a wide range of programs from 2017 to 2025, aiming at both real-world implementation and use-inspired research. These funding documents were provided by government agencies (11 funding documents), philanthropic foundations (10), nonprofit organizations (7), and private technology companies (8) based in English-speaking countries. These funders were headquartered in the United States (26), the United Kingdom (5), Canada (4), Sweden (3), Germany (1), Ireland (1), and Australia (1). These funding programs targeted applications from countries including the United States (5), the United Kingdom (1), Ireland (1), and South Africa (1), as well as broader regions including globally (14), Africa (9), Latin America (2), and Low- and Medium-Income Countries (2).
@@ 3.2. Thematic Analysis on Funding Documents
In the third coding stage, the first author coded all the funding documents, with our research questions in mind. Each funding document’s codes were reviewed by another team member. The first author then resolved any (though few) disagreements (denoted as comments on Google documents) with each code reviewer through online discussions. These discussions involved clarifying the meaning of the first author’s original codes and discussing whether they aligned with the original texts in the funding documents. We did not calculate intercoder reliability, since reflexive thematic analysis treats coding as an interpretive process rather than one that seeks positivist (quantitative) reliability measures (Braun and Clarke, 2019). All authors then conducted iterative affinity diagramming on an online whiteboard and discussed emerging high-level themes from the codes. The purpose of affinity diagramming was to categorize low-level codes (at the sentence level) and surface higher-level themes across the funding documents (at the section level and beyond), as well as find connections between these themes. We collaboratively characterized these high-level themes as a spectrum between a more techno-centric and a more balanced approach, using the lens of the prior literature synthesized in Section 2.2.333Appendix B provides screenshots of our coding process and affinity diagrams.